# Deep learning approach for identification of Hii regions during reionization in 21-cm observations – III. image recovery

### v1

@@ Abstract
pubyear: 2024pagerange: Deep learning approach for identification of Hii regions during reionization in 21-cm observations – III. image recovery14
@@ 1 Introduction
The 21-cm signal from the Epoch of Reionization (EoR) provides a unique window into the early Universe, revealing the processes of cosmic structure formation and the distribution of ionizing sources that transitioned the intergalactic medium (IGM) of our Universe from an initial cold and neutral to a final hot and ionized state (e.g. Zaldarriaga et al., 2004; Furlanetto et al., 2006; Pritchard & Loeb, 2012). Several ongoing radio experiments, such as Low-frequency Array (LOFAR; e.g. Mertens et al., 2020; Ghara et al., 2020), Murchison Widefield Array (MWA; e.g. Tingay et al., 2013; Trott et al., 2020) and Hydrogen Epoch of Reionization Array (HERA; e.g. The HERA Collaboration et al., 2022; HERA Collaboration et al., 2023). The upcoming Square Kilometre Array [^1] https://www.skao.int/en Low-Frequency [^2] https://www.skao.int/en/explore/telescopes/ska-low component (SKA-Low) is expected to enhance our ability to observe this signal significantly, offering unprecedented sensitivity and spatial resolution (e.g. Koopmans et al., 2015). However, extracting meaningful information from these observations presents substantial challenges, primarily due to the overwhelming foreground contamination and the inherent non-Gaussianity of the 21-cm signal.
@@
The foreground signal consists of galactic and extra-galactic components with various shapes, sizes and flux densities of the point and extended sources. At low radio frequencies, the foreground removal is a major obstacle, as the flux densities of this signal often far exceed the 21-cm signal of several orders of magnitude (e.g. Ali et al., 2008; Jelić et al., 2008, 2010; Ghosh et al., 2012). Traditional foreground subtraction techniques struggle to manage the complex, diffuse nature of these foregrounds, leading to residuals that can obscure or distort the signal of interest. To address this, advanced methods are required to effectively isolate and recover the 21-cm signal from the observed data (e.g. Harker et al., 2009, 2010; Liu & Tegmark, 2011; Liu et al., 2014a, b; Chapman et al., 2012, 2016; Pober et al., 2016; Mertens et al., 2018; Hothi et al., 2021; Kern & Liu, 2021; Acharya et al., 2024).
@@
In recent years, machine learning techniques have been consolidating their role as one of the most widely used tools employed by the cosmological and astrophysical scientific community for data analysis (see e.g. Dvorkin et al., 2022; Lahav, 2023, for on overview). In particular, U-Nets are becoming the standard in studies with tasks that require processing images or 3D data with structural correlations (e.g. He et al., 2015; Münchmeyer & Smith, 2019; Perraudin et al., 2019; Gupta & Reichardt, 2020; Jeffrey et al., 2020; Guzman & Meyers, 2021; Gupta & Reichardt, 2021; Makinen et al., 2021a; Bretonnière et al., 2021; Dong et al., 2021; Montefalcone et al., 2021; Han et al., 2021; Jeffrey et al., 2021; Casas et al., 2022; Li et al., 2022; Ostdiek et al., 2022a, b; Schaurecker et al., 2022; Piras et al., 2023). In recent years, the field of 21-cm cosmology has employed U-shaped networks (U-Nets) to tackle problems such as foreground contamination or performing other complex non-linear tasks (Villanueva-Domingo & Villaescusa-Navarro, 2021; Gagnon-Hartman et al., 2021; Bianco et al., 2021; Masipa et al., 2023; Bianco et al., 2024; Kennedy et al., 2024; Shi et al., 2024; Sabti et al., 2024, e.g.). Although experiencing rapid success, the full potential of these networks still remains uncovered. Recently, Chen et al. (2024) investigated the strengths and limitations of this technique.
@@
In our previous work, we focused on the detection of neutral regions with a segmentation network known as SegU-Net and tested this network’s capability in the presence of foreground residuals (Bianco et al., 2024, hereafter Paper II). In this paper, we want to take a step further and develop a network that also recovers the 21-cm signal. For this purpose, we go beyond the conventional development of U-Net and we propose a new architecture motivated by work from a completely different field. The authors of Tofighi et al. (2019) and El Jurdi et al. (2020) proposed a modified version of U-Net that includes an additional input to enhance their segmentation detection. We took inspiration from their work and decided to develop a new network with this alternative architecture to enhance the recovery of the 21-cm signal from foreground contamination images from SKA-Low tomographic images. Therefore, we constructed RecU-Net for recovering the signal. Apart from the standard elements of a U-Net (Ronneberger et al., 2015) architecture (encode, decoder and skip layers), we construct an added component to the architecture that we named the Interception Convolutional Block. A series of convolution and pooling layers process an additional input that we refer to as the prior maps, and the result intercepts the skipping layer of the U-Net to enforce prior information onto the network encoder and decoder.
@@
We perform a complete study assessing the new architecture and the advantage of including different binary prior maps in recovering 21-cm signals from foreground-contaminated EoR data. We demonstrate that by using a binary field derived from the volume-averaged neutral fraction field $x_{\mathrm{HI}}(\mathbf{r},z)$, the SERENEt frameworks achieves more stable predictions and a better 21-cm signal recovery. Furthermore, we demonstrate that any meaningful prior information could be used without requiring retraining the network. Finally, we show that SERENEt could potentially be employed for synergy observations between SKA-Low and high redshift galaxy observations, such as the James Webb Space Telescope (JWST), Euclid and Nancy Grace Roman Space Telescope. This synergy would be potentially beneficial to both experiments by enhancing 21-cm recovery and understanding the relation between the early sources of reionization and properties of the IGM (e.g. Zackrisson et al., 2020).
@@
Figure 1: Schematic diagram of the SERENEt pipeline. The first step consists of pre-processing the observed image, $I_{\mathrm{obs}}$, to remove the largest contribution of the foreground contamination resulting in a residual image, $I_{\mathrm{res}}$. The second step involves identifying neutral regions with the SegU-Net architecture. The segmentation provides a binary map, $I_{\mathrm{B}}$, that, combined with the residual image, provides the input of the third step, which consists of recovering the 21-cm signal with the RecU-Net architecture. The combination of the two networks constitutes SERENEt that constructs the final image.
@@ 2.1 Modelling structure formation and reionization
We employ a semi-numerical code, named 21cmFAST (e.g. Mesinger et al., 2011; Murray et al., 2020) that allows simulating both the cosmic structure formation and ionization of the IGM. The large-scale dark matter distribution at any cosmic time is determined using second-order Lagrangian Perturbation theory (2LPT; Scoccimarro et al., 1998) in simulations volumes on grids of $128^{3}$ and length of 256 Mpc along each direction. While several more accurate modelling frameworks, such as pyC2Ray (Hirling et al., 2024) and CRASH (Maselli et al., 2003), exist, we chose this semi-numerical code as it is computationally inexpensive enough to allow us to create a large dataset for training our deep learning model SERENEt.
@@
where $\zeta$ is the ionizing source efficiency and $f_{\mathrm{coll}}$ is the collapse fraction of dark matter haloes of mass more than $M_{\mathrm{min}}$. This quantity is estimated at this cell assuming a spherical top-hat filter of radius $R$. Several $R$ values are tested until a maximum value of $R_{\mathrm{max}}$.
@@ 2.2 Modelling the observational effects
We constructed synthetic observations of SKA-Low by simulating the uv tracks using the planned antenna locations [^3] These locations were obtained from the available SKAO document: d18-SKA-TEL-SKO-0000422_02_SKA1_LowConfigurationCoordinates-1.pdf, which was prepared by Peter E. Dewdney and Robert Braun.. The observations assume a total observation time of 1000 hours, with data collected for 6 hours each day. This extended observation time helps to reduce instrumental noise sufficiently to resolve meaningful features in 21-cm signal images (Giri et al., 2018b). Additionally, we consider the antenna distribution contained within a maximum diameter of 2 km around the core of SKA-Low that reduced the resolution of the 21-cm signal image data (Mellema et al., 2015; Giri et al., 2018a).
@@ 3 Machine Learning Application
Figure 2: Architecture of the RecU-Net model. The standard U-shape is modified by introducing Interception Convolutional Blocks, which consist of a series of pooling and convolutional operations that intercept and combine with the output of the skip connections. We use binary images from SegU-Net predictions in this case, but any data can be used as prior in this architecture.
@@ 3.1 SERENEt pipeline
Several previous works have demonstrated that foreground mitigation can be improved by combining machine learning techniques with standard mitigation approaches such as principal component analysis (PCA) and polynomial fitting (e.g. Li et al., 2019; Makinen et al., 2021b). Therefore, our pipeline follows this approach and is divided into three main steps. In Figure 1, we show a schematic diagram of this pipeline. In the first step of our pipeline, we pre-process the foreground-contaminated image, $I_{\rm obs}$, of mesh-size $128^{2}$, with a simple PCA method [^4] We employ the module implemented in the SciPy software package that can be found at https://docs.scipy.org.. This method models the foreground by applying PCA at each pixel along the frequency direction. As the foreground signal is expected to be smooth along frequency, the components with the largest contributions should contain the foreground signal. The choice of the number of PCA modes corresponding to the foreground signal affects the level of residuals (e.g. Harker et al., 2009). In this study, we remove the first four modes from a PCA decomposition in the tomographic dataset and reconstruct the image, $I_{\rm res}$, with the remaining components. This step allows us to remove the significant contribution of the foreground signal. However, as discussed next, the neural network will handle systematic noise and the remaining residual foreground signal.
@@ 3.3 Metrics for Training and Validation
where $TP=\sum_{i}y_{i}\hat{y}_{i}$ indicates the true positive, $TN=\sum_{i}(1-y_{i})(1-\hat{y}_{i})$ is the true negative, $FP=\sum_{i}y_{i}(1-\hat{y}_{i})$ is the false positive, and $FN=\sum_{i}(1-y_{i})\hat{y}_{i}$ is the false negative. $1$ indicates a neutral region in our binary maps, while 00 is ionized. Therefore, in our case, $TP$ ($TN$) indicates the number of pixels in the image that are correctly identified as neutral (ionized). Instead, $FP$ and $FN$ quantifies the corresponding wrong classification. This metric ranges between $-1\leq r_{\phi}\leq 1$; a positive value indicates a correlation between the predicted binary map and the ground truth image. A negative value instead indicates anti-correlation, while values close to zeros indicate a completely random classification.
@@ 4 Results
In this section, we analyze the prediction of our networks. In particular, we consider three cases. In the first case, we consider 21-cm images reconstructed by a simple U-Net. This is equivalent to employing RecU-Net, but without the Interception Convolutional Block; in this case, we want to test if the modified architecture of RecU-Net provides an actual advantage. We refer to this setup as the no prior test. We consider the modified architecture in the other two cases, as illustrated in §3.2. The former considers the perfect scenario where the binary prediction, thus the prior, from SegU-Net obtains a perfect score. We refer to this setup as the perfect prior test, and in this case, we employ the actual ground truth binary maps in the training and validation of RecU-Net. Our last test is the realistic prior setup. Here, we employ the trained SegU-Net as presented in Paper II and provide a realistic scenario where the recovery of the 21-cm image by RecU-Net will also depend on the accuracy of the binary prior predicted by SegU-Net.
@@ 4.2 Effect of the Prior on the 21-cm Signal Predictions
The first observation is that while the simple U-shaped architecture without binary prior maps recovers the 21-cm signal with satisfactory accuracy ($0.850.65<R^{2}<$ 0.85), it underperforms compared to cases that utilize the prior $I_{\mathrm{B}}$. As expected, the perfect prior case yields the best predictions, with $R^{2}>0.9$ for the three reionization stages presented in Figure 4. In contrast, the realistic prior provides results that fall between the other two cases.
@@
Figure 6: Statistical results of SERENEt predictions for the realistic prior case on the random dataset. Left panel: correlation between the $R^{2}$ score and the true volume-averaged neutral fraction $\overline{x}_{\mathrm{HI,true}}$. Each point represents 21-cm signal data at the corresponding redshift, indicated by the colour bar. Right panel: correlation between the $R^{2}$ score and the binary score $r\phi$ of the realistic prior predicted by SegU-Net. The colour indicates the $\overline{x}_{\mathrm{HI,true}}$ of the image. Ideally, all points should be close to the top right corner, representing a perfect full pipeline. However, we observe that data with a high $\overline{x}_{\mathrm{HI,true}}$ has good SegU-Net predictions but less accurate SERENEt results, while data with a low $\overline{x}_{\mathrm{HI,true}}$ shows poor SegU-Net predictions but accurate SERENEt results. Both panels show 68 and 95 per cent confidence intervals with green solid and dashed contours.
@@
The advantage of the prior is particularly evident from the redshift evolution of the $R^{2}$ score, shown in Figure 5. The $R^{2}$ score for the no prior case (blue line) is substantially lower than that for the other tests—approximately $5-35$ per cent below the perfect prior (orange line) and up to $25$ per cent below the realistic prior (green line). Furthermore, the accuracy of the perfect prior scenario remains constant at $R^{2}\approx 0.97$ until $z\simeq 9$, when $\overline{x}_{\mathrm{HI}}\gtrsim 0.7$. In contrast, the accuracy in the other cases gradually decreases at higher redshifts.

### v2

@@ Abstract
pubyear: 2024pagerange: Deep learning approach for identification of Hii regions during reionization in 21-cm observations – III. image recoveryB
@@ 1 Introduction
The 21-cm signal from the Epoch of Reionization (EoR) provides a unique window into the early Universe, revealing the processes of cosmic structure formation and the distribution of ionizing sources that transitioned the intergalactic medium (IGM) of our Universe from an initial cold and neutral to a final hot and ionized state (e.g. Zaldarriaga et al., 2004; Furlanetto et al., 2006; Pritchard & Loeb, 2012). Several ongoing radio experiments, such as the Low-frequency Array (LOFAR; e.g. Mertens et al., 2020; Ghara et al., 2020), Murchison Widefield Array (MWA; e.g. Tingay et al., 2013; Trott et al., 2020) and Hydrogen Epoch of Reionization Array (HERA; e.g. The HERA Collaboration et al., 2022; HERA Collaboration et al., 2023) aim to observe this transition in our Universe. The upcoming Square Kilometre Array [^1] https://www.skao.int/en Low-Frequency [^2] https://www.skao.int/en/explore/telescopes/ska-low component (SKA-Low) is expected to enhance our ability to observe this signal significantly, offering unprecedented sensitivity and spatial resolution (e.g. Koopmans et al., 2015). However, extracting meaningful information from these observations presents substantial challenges, primarily due to the overwhelming foreground contamination and the inherent non-Gaussianity of the 21-cm signal.
@@
The foreground signal consists of galactic and extra-galactic components with various shapes, sizes and flux densities of the point and extended sources. At low radio frequencies, the foreground removal is a major obstacle, as the flux densities of this signal often far exceed the 21-cm signal by several orders of magnitude (e.g. Ali et al., 2008; Jelić et al., 2008, 2010; Ghosh et al., 2012). Traditional foreground subtraction techniques struggle to manage the complex, diffuse nature of these foregrounds, leading to residuals that can obscure or distort the signal of interest. To address this, advanced methods are required to effectively isolate and recover the 21-cm signal from the observed data (e.g. Harker et al., 2009, 2010; Liu & Tegmark, 2011; Liu et al., 2014a, b; Chapman et al., 2012, 2016; Pober et al., 2016; Mertens et al., 2018; Hothi et al., 2021; Kern & Liu, 2021; Acharya et al., 2024).
@@
In recent years, machine learning techniques have been consolidating their role as one of the most widely used tools employed by the cosmological and astrophysical scientific community for data analysis (see e.g. Dvorkin et al., 2022; Lahav, 2023, for on overview). In particular, U-Nets are becoming the standard in studies with tasks that require processing images or 3D data with structural correlations (e.g. He et al., 2015; Münchmeyer & Smith, 2019; Perraudin et al., 2019; Gupta & Reichardt, 2020; Jeffrey et al., 2020; Guzman & Meyers, 2021; Gupta & Reichardt, 2021; Makinen et al., 2021a; Bretonnière et al., 2021; Dong et al., 2021; Montefalcone et al., 2021; Han et al., 2021; Jeffrey et al., 2021; Casas et al., 2022; Li et al., 2022; Ostdiek et al., 2022a, b; Schaurecker et al., 2022; Piras et al., 2023). In recent years, the field of 21-cm cosmology has employed U-shaped networks (U-Nets) to tackle problems such as foreground contamination or performing other complex non-linear tasks (Villanueva-Domingo & Villaescusa-Navarro, 2021; Gagnon-Hartman et al., 2021a; Bianco et al., 2021; Masipa et al., 2023; Hiegel et al., 2023; Bianco et al., 2024; Kennedy et al., 2024; Shi et al., 2024; Sabti et al., 2024, e.g.). Although experiencing rapid success, the full potential of these networks still remains uncovered. Recently, Chen et al. (2024) investigated the strengths and limitations of this technique.
@@
In our previous work, we focused on the detection of neutral regions with a segmentation network known as SegU-Net and tested this network’s capability in the presence of foreground residuals (Bianco et al., 2024, hereafter Paper II). In this paper, we want to take a step further and develop a network that also recovers the 21-cm signal. For this purpose, we go beyond the conventional development of U-Net and we propose a new architecture motivated by work from a completely different field. The authors of Tofighi et al. (2019) and El Jurdi et al. (2020) proposed a modified version of U-Net that includes an additional input to enhance their segmentation detection. We took inspiration from their work and decided to develop a new network with this alternative architecture to enhance the recovery of the 21-cm signal from foreground contamination images from SKA-Low tomographic images. Therefore, we constructed RecU-Net for recovering the signal. Apart from the standard elements of a U-Net (Ronneberger et al., 2015) architecture (encode, decoder and skip layers), we construct an added component to the architecture that we named the Interception Convolutional Block. A series of convolution and pooling layers processes an additional input that we refer to as the prior maps, and the result intercepts the skipping layer of the U-Net to enforce prior information onto the network encoder and decoder.
@@
We perform a complete study assessing the new architecture and the advantage of including different binary prior maps in recovering 21-cm signals from foreground-contaminated EoR data. We demonstrate that by using a binary field derived from the volume-averaged neutral fraction field $x_{\mathrm{HI}}(\mathbf{r},z)$, the SERENEt frameworks achieve more stable predictions and a better 21-cm signal recovery. Furthermore, we demonstrate that any meaningful prior information could be used without requiring retraining the network. Finally, we show that SERENEt could potentially be employed for synergy observations between SKA-Low and high redshift galaxy observations, such as the James Webb Space Telescope (JWST), Euclid and Nancy Grace Roman Space Telescope. This synergy would be potentially beneficial to both experiments by enhancing 21-cm recovery and understanding the relation between the early sources of reionization and properties of the IGM (e.g. Zackrisson et al., 2020).
@@
Figure 1: Schematic diagram of the SERENEt pipeline. The first step consists of pre-processing the observed image, $I_{\mathrm{obs}}$, to remove the largest contribution of the foreground contamination, resulting in a residual image, $I_{\mathrm{res}}$. The second step involves identifying neutral regions with the SegU-Net architecture. The segmentation provides a binary map, $I_{\mathrm{B}}$, that, combined with the residual image, provides the input of the third step, which consists of recovering the 21-cm signal with the RecU-Net architecture. The combination of the two networks constitutes SERENEt that constructs the final image.
@@ 2.1 Modelling structure formation and reionization
We employ a semi-numerical code, named 21cmFAST (e.g. Mesinger et al., 2011; Murray et al., 2020) that allows simulating both the cosmic structure formation and ionization of the IGM. The large-scale dark matter distribution at any cosmic time is determined using second-order Lagrangian Perturbation theory (2LPT; Scoccimarro et al., 1998) in simulation volumes on grids of $128^{3}$ and length of 256 Mpc along each direction. While several more accurate modelling frameworks, such as pyC2Ray (Hirling et al., 2024) and CRASH (Maselli et al., 2003), exist, we chose this semi-numerical code as it is computationally inexpensive enough to allow us to create a large dataset for training our deep learning model SERENEt.
@@
where $\zeta$ is the ionizing source efficiency and $f_{\mathrm{coll}}$ is the collapse fraction of dark matter haloes of mass more than $M_{\mathrm{min}}$. This quantity is estimated at this cell assuming a spherical top-hat filter of radius $R_{\mathrm{s}}$. Several $R_{\mathrm{s}}$ values are tested until a maximum value of $R_{\mathrm{mfp}}$.
@@ 2.2 Modelling the observational effects
We constructed synthetic observations of SKA-Low by simulating the uv tracks using the planned antenna locations, conforming to the SKAO configuration coordinates plan (Dewdney, 2016). The observations assume a total observation time of 1000 hours, with data collected for 6 hours each day. This extended observation time helps to reduce instrumental noise sufficiently to resolve meaningful features in 21-cm signal images (Giri et al., 2018b). Additionally, we consider the antenna distribution contained within a maximum diameter of 2 km around the core of SKA-Low that reduced the resolution of the 21-cm signal image data (Mellema et al., 2015; Giri et al., 2018a).
@@ 3 Machine Learning Application
Figure 2: Architecture of the RecU-Net model. The standard U-shape is modified by introducing Interception Convolutional Blocks, which consist of a series of pooling and convolutional operations that intercept and combine with the output of the skip connections. In this case, we use binary images from SegU-Net predictions, but any data can be used as prior in this architecture.
@@ 3.1 SERENEt pipeline
Several previous works have demonstrated that foreground mitigation can be improved by combining machine learning techniques with standard mitigation approaches such as principal component analysis (PCA) and polynomial fitting (e.g. Li et al., 2019; Makinen et al., 2021b). Therefore, our pipeline follows this approach and is divided into three main steps. In Figure 1, we show a schematic diagram of this pipeline. In the first step of our pipeline, we pre-process the foreground-contaminated image, $I_{\rm obs}$, of mesh-size $128^{2}$, with a simple PCA method [^3] We employ the module implemented in the SciPy software package that can be found at https://docs.scipy.org.. This method models the foreground by applying PCA at each pixel along the frequency direction. As the foreground signal is expected to be smooth along frequency, the components with the most significant contributions should contain the foreground signal. The choice of the number of PCA modes corresponding to the foreground signal affects the level of residuals (e.g. Harker et al., 2009). This study removes the first four modes from a PCA decomposition in the tomographic dataset and reconstructs the image, $I_{\rm res}$, with the remaining components. This step allows us to remove the significant contribution of the foreground signal. However, as discussed next, the neural network will handle systematic noise and the remaining residual foreground signal.
@@ 3.3 Metrics for Training and Validation
where $TP=\sum_{i}y_{i}\hat{y}_{i}$ indicates the true positive, $TN=\sum_{i}(1-y_{i})(1-\hat{y}_{i})$ is the true negative, $FP=\sum_{i}y_{i}(1-\hat{y}_{i})$ is the false positive, and $FN=\sum_{i}(1-y_{i})\hat{y}_{i}$ is the false negative. $1$ indicates a neutral region in our binary maps, while 0 is ionized. Therefore, in our case, TPTP (TNTN) indicates the number of pixels in the image that are correctly identified as neutral (ionized). Instead, $FP$ and $FN$ quantifies the corresponding wrong classification. This metric ranges between $1\leq r_{\phi}\leq 1$; a positive value indicates a correlation between the predicted binary map and the ground truth image. A negative value instead indicates anti-correlation, while values close to zeros indicate a completely random classification.
@@ 4 Results
In this section, we analyse the prediction of our networks. In particular, we consider three cases. In the first case, we consider 21-cm images reconstructed by a simple U-Net. This is equivalent to employing RecU-Net, but without the Interception Convolutional Block; in this case, we want to test if the modified architecture of RecU-Net provides an actual advantage. We refer to this setup as the no prior test. We consider the modified architecture in the other two cases, as illustrated in §3.2. The former considers the perfect scenario where the binary prediction, thus the prior, from SegU-Net obtains a perfect score. We refer to this setup as the perfect prior test, and in this case, we employ the actual ground truth binary maps in the training and validation of RecU-Net. Our last test is the realistic prior setup. Here, we employ the trained SegU-Net as presented in Paper II and provide a realistic scenario where the recovery of the 21-cm image by RecU-Net will also depend on the accuracy of the binary prior predicted by SegU-Net.
@@ 4.2 Effect of the Prior on the 21-cm Signal Predictions
The first observation is that while the simple U-shaped architecture without binary prior maps recovers the 21-cm signal with satisfactory accuracy ($0.65<R^{2}<0.85$), it underperforms compared to cases that utilize the prior $I_{\mathrm{B}}$. As expected, the perfect prior case yields the best predictions, with $R^{2}>0.9$ for the three reionization stages presented in Figure 4. In contrast, the realistic prior provides results that fall between the other two cases.
@@
Figure 6: Statistical results of SERENEt predictions for the realistic prior case on the random dataset. Left panel: correlation between the $R^{2}$ score and the true volume-averaged neutral fraction $\overline{x}_{\mathrm{HI,true}}$. Each point represents 21-cm signal data at the corresponding redshift, indicated by the colour bar. Right panel: correlation between the $R^{2}$ score and the binary score $r_{\phi}$ of the realistic prior predicted by SegU-Net. The colour indicates the $\overline{x}_{\mathrm{HI,true}}$ of the image. Ideally, all points should be close to the top right corner, representing a perfect full pipeline. However, we observe that data with a high $\overline{x}_{\mathrm{HI,true}}$ has good SegU-Net predictions but less accurate SERENEt results, while data with a low $\overline{x}_{\mathrm{HI,true}}$ shows poor SegU-Net predictions but accurate SERENEt results. Both panels show 68 and 95 per cent confidence intervals with green solid and dashed contours.
@@
The advantage of the prior is particularly evident from the redshift evolution of the $R^{2}$ score, shown in Figure 5. The $R^{2}$ score for the no prior case (blue line) is substantially lower than that for the other tests—approximately 5355-35 per cent below the perfect prior (orange line) and up to $25$ per cent below the realistic prior (green line). Furthermore, the accuracy of the perfect prior scenario remains constant at $R^{2}\approx 0.97$ until $z\simeq 9$, when $\overline{x}_{\mathrm{HI}}\gtrsim 0.7$. In contrast, the accuracy in the other cases gradually decreases at higher redshifts. This trend depends on the increasing noise level, but mainly on the number of components we remove in the PCA pre-processing. Removing more components allows for a better mitigation of the foreground at higher redshift, lightening the accuracy decrease at higher redshift but increasing the portion of the 21-cm signal removed at lower redshifts. Following the approach in Paper II, we removed the first four components, which contained the most significant contribution of the foreground contamination, and as it provides a balanced accuracy between high and low redshift.
Moreover, in Figure 5, one can notice a sudden drop in accuracy at different redshifts in all the scenarios, in the fiducial model, the most evident one occurs at $z\sim 10$. These events are due to numerical artifacts in the imperfect PCA foreground decomposition, as we can observe them in the perfect prior only at high redshift, $z>9$. We statistically explore these trends further in the next section, §4.3.